{F} Determine the length of the parametric curve given by the following parametric equations. x=3sin(t) y=3cos(t) 0<t<2π


1
Expert's answer
2022-02-01T07:39:00-0500
"dx/dt=3\cos(t), dy/dt=-3\sin(t)"

"L=\displaystyle\int_{0}^{2\pi}\sqrt{(dx/dt)^2+(dy/dt)^2}dt"

"=\displaystyle\int_{0}^{2\pi}\sqrt{(3\cos(t))^2+(-3\sin(t))^2}dt"

"=3\displaystyle\int_{0}^{2\pi}dt=3[t]\begin{matrix}
 2\pi \\
 0
\end{matrix}=3(2\pi-0)=6\pi(units)"

"L=6\pi\ units"